Calculating Physician Service Statistics

You are on the Chargemaster Committee at Pike's Peak Clinic. An annual review of services and charges is being conducted for necessary revisions. Table 8 shows some statistics reported on the physician's services at Pike's Peak Clinic last Tuesday.

Table 8 Pike's Peak Clinic Physician Services Statistics for Tuesday

Pike’s Peak Clinic
Physician Service Statistics for Tuesday

Physician         Service A                        Service B Service C

Truba                     14                                 10                                     18               

Wooley                     9                             14 22   

Howe                        6                          18                                        5    

Masters                   7                              12                                      20   

It takes twice as long to perform Service C, so the physicians are proposing that Service C should count as two services for the purpose of calculating workload.   

If service C counts twice as much as Service A or Service B, which physician provided the most services on Tuesday?
Answer #1

From the given data, the following Table is calculated:

Physician Service A Service B Service C Total number of services
Truba 14 10 18 14 + 10 + (18 X 2) = 60
Wooley 9 14 22 9 + 14 + (22 X 2) = 67
Howe 6 18 5 6 + 18 + (5 X 2) = 34
Masters 7 12 20 7 + 12 + (20 X 2) = 59

So

The Physician Wooley provided the most services on Tuesday because Total number of services of Wooley = 67 is the maximum of the 4 physicians.